A rectangle has odd dimensions

Grade:11

1 Answers

Komal
8 years ago
Dimensions: 2n-1, 2n+1
Perimeter: 2 (2n -1 + 2n+1) = 2 (4n) = 8n.

So we need 8n > 35, or n > 35/8 = 4.375. But n must be an integer, so n=5 and the dimensions are 9 and 11.